# HoloDx: Knowledge- and Data-Driven Multimodal Diagnosis of Alzheimer’s Disease

### v1

@@ Abstract
Accurate diagnosis of Alzheimer’s disease (AD) requires effectively integrating multimodal data and clinical expertise. However, existing methods often struggle to fully utilize multimodal information and lack structured mechanisms to incorporate dynamic domain knowledge. To address these limitations, we propose HoloDx, a knowledge- and data-driven framework that enhances AD diagnosis by aligning domain knowledge with multimodal clinical data. HoloDx incorporates a knowledge injection module with a knowledge-aware gated cross-attention, allowing the model to dynamically integrate domain-specific insights from both large language models (LLMs) and clinical expertise. Also, a memory injection module with a designed prototypical memory attention enables the model to retain and retrieve subject-specific information, ensuring consistency in decision-making. By jointly leveraging these mechanisms, HoloDx enhances interpretability, improves robustness, and effectively aligns prior knowledge with current subject data. Evaluations on five AD datasets demonstrate that HoloDx outperforms state-of-the-art methods, achieving superior diagnostic accuracy and strong generalization across diverse cohorts. The source code will be released upon publication acceptance.
@@ 1 Introduction
Alzheimer’s disease (AD) is a progressive neurodegenerative disorder that leads to memory loss, cognitive decline, and ultimately, loss of independence. Accurate diagnosis of AD is crucial for implementing appropriate treatment strategies, planning for future care, and providing support to patients and their families. Recent advancements [1, 2] in passive immunotherapies, such as Lecanemab and Donanemab, have shown promise in targeting amyloid-beta plaques, a hallmark of AD pathology. However, diagnosing AD, especially in its early stages, remains challenging due to subtle clinical symptoms and overlap with normal aging processes. Traditional diagnostic approaches, often relying on unimodal data such as neuroimaging or clinical assessments, struggle to capture these nuanced changes [3, 4]. To address these limitations, multimodal learning has gained attention for its ability to integrate diverse data sources, such as neuroimaging, clinical records, and genetic information, offering a more holistic perspective on AD’s early pathological changes [5, 6]. Despite the promise of data-driven approaches, significant challenges remain: 1) Data bias, which focuses on prevalent patterns while often neglecting rare and atypical cases, and 2) A lack of interpretability, as many models function as opaque black boxes, making it difficult to trace causal relationships and ensure reliability. These limitations hinder the consistency, robustness, and trustworthiness of existing computer-aided diagnosis (CAD) systems. Addressing these shortcomings requires a synergistic paradigm, as shown in Fig. 1, which integrates the strengths of multimodal data-driven models with domain knowledge, enhancing interpretability, reliability, and generalizability in the context of AD and its early diagnosis.
@@ 2.1 Multimodal CAD
The integration of diverse modalities of medical data has become a prominent approach in advancing medical diagnosis [10, 11], enabling machine learning models to identify disease patterns more effectively than relying on single modalities. For example, Kim et al. [12] develop a heterogeneous graph learning method to fuse multimodal medical data efficiently. Similarly, Zhou et al. [6] introduce the IRENE model, which employs Transformer-based architectures to combine modality-specific low-level embeddings for diagnostic purposes. Multimodal integration has also succeeded in diagnosing conditions such as breast cancer, where combining imaging and genomic information has improved diagnostic outcomes [13]. Chen et al. [5] propose a transformer-based framework for aligning and fusing multimodal medical data to classify Alzheimer’s disease. These efforts highlight the growing potential of multimodal CAD systems to improve diagnostic precision and reliability. However, these approaches often rely solely on data-driven techniques, which limits their ability to handle complex medical scenarios where domain expertise is crucial. Without integrating domain knowledge, such models may fail to resolve conflicts between modalities or interpret rare and ambiguous cases effectively. This gap underscores the need for combining multimodal data integration with domain knowledge to enhance interpretability, consistency, and diagnostic reliability.
@@ 2.2 Knowledge-driven Multimodal CAD
Our approach bridges this gap by dynamically enriching the knowledge bank with general medical insights derived from large language models (LLMs) and specific knowledge contributed by domain experts, tailored to the factors associated with each subject. This ensures that multimodal data is augmented with both broad and context-specific representations, enabling our method to provide more precise, interpretable, and robust predictions, effectively addressing key limitations in existing knowledge-driven frameworks.
@@ 3.1.1 Factor Extraction
Expert: The increase of p-tau217 in the serum of AD patients is usually early and significant, and its diagnostic power is generally superior to that of Abeta42/40 and ptau181. Therefore, p-tau217 holds high clinical value in the early diagnosis of Alzheimer’s disease. | |
@@ 3.1.2 Factor Knowledge Generation
Here, “factor” refers to a specific factor , as listed in Table 1. Examples include the Mini-Mental State Examination (MMSE) scores collected from cognitive tests. The generated responses offer comprehensive descriptions of each factor and its relevance to diagnosing AD, as presented in the LLM knowledge column of Table 2.
@@ 3.2 Multimodal and Knowledge Feature Extraction
Unimodal Encoder. We employ unimodal encoders to process both visual inputs and textual inputs . For the visual modality, a 3D Vision Transformer [4] is used to encode the input images into vector representations, generating an image embedding that includes a [CLS] token to represent the global image feature. For the textual modality, we use a Longformer Transformer [19], which maps long text representations into feature embeddings. Similarly, a [CLS] token is appended to the beginning of the text input to summarize the whole sentence representation. Each unimodal encoder consists of transformer blocks comprising self-attention (SA) layers and feed-forward networks (FFN).
@@ 3.3 Knowledge Injection Module
These multimodal embeddings first undergo self-attention (SA) to model interdependencies across modalities:
@@
The full layer transformation integrates these components through residual learning:
@@
The final patient representation aggregates modality-specific information through concatenation of [CLS] tokens:
where , , and correspond to the [CLS] tokens of visual, textual, and knowledge embeddings respectively.
@@ 3.5.1 Alignment
We employ the image-text contrastive (ITC) loss [23] to align image features and text features generated by the image and text encoders. The ITC loss maximizes similarity for positive image-text pairs while suppressing negative pairs, implemented through normalized cross-entropy over all pairwise similarities. To enhance consistency between input data and domain knowledge in disease diagnosis, we propose the Knowledge-Data Contrastive loss (KDC). This mechanism maintains two dynamically updated queues: a Data queue storing concatenated multimodal features , and a Knowledge queue containing domain embeddings generated by a dedicated knowledge encoder. The KDC loss follows the contrastive paradigm of , pulling matched knowledge-data pairs closer while distancing mismatched pairs. Momentum encoders [24] continuously refresh both queues to maintain feature relevance.
@@ 4.1 Datasets and Settings
In accordance with established clinical standards [35, 36, 37], we validate our model on two classification tasks. The first task distinguishes between NC and Cognitive Impairment (CI), where CI includes both MCI and AD. The second task differentiates between NC and Early Cognitive Impairment (ECI), with ECI being used synonymously with MCI, a crucial stage for the preclinical identification of AD. We split the ADNI, AIBL, and RENJI datasets subject-wise into training, validation, and test sets with proportions of 70%, 10%, and 20%, respectively. Also, we perform zero-shot tests on OASIS and MIRIAD datasets to assess inter-center generalization.
@@
Baselines. To evaluate the performance of our model, we compare it with three groups of baselines: (1) Image-only group. This group includes two recent methods commonly used in image classification: 3D ResNet [26], and 3D ViT [4]. (2) Text-only group. We select three prominent text-only baselines: BERT [27], Roberta [28], and Longformer [19]. (3) Multimodal group. This group includes three recent transformer-based models that fuse multimodal information for classification: GIT [29], IRENE [6], and Alifuse [5].
@@
Implementation details For pre-processing, each image is resized and cropped to , with a spacing of [1.0,1.0,1.0] and an image patch size of . The maximum input text length of textual input is set to 512. Following common settings in [25], we set all hidden feature sizes to 768 and the number of heads in multi-head attention (MHA) to 12. The number of layers in the unimodal encoder is set to 12, while the number of layers in the decoder is set to 6. The number of layers in the knowledge and the memory modules is set to 6. The iteration T is 100. The values for are empirically set to 1, 1, and 1, respectively. Our model is trained on four NVIDIA RTX 3090 GPUs with a batch size of , using the AdamW optimizer with a learning rate of 2e-5. Evaluation metrics include diagnosis accuracy (ACC), Area Under Curve (AUC), sensitivity (SEN), and specificity (SPE).
@@ 4.2.3 Ablation Studies
(1) Impact of loss terms. We evaluate the contribution of each loss, i.e., the classification loss , two contrastive losses and , and the combined restorative loss for images and texts. Incorporating all loss components is optimal for maximizing performance across all metrics, as reported in Table 7(a). Using only , the model performance experiences a reduction of 2.04% in ACC and 2.01% in AUC compared to the best-performing configuration. Adding and alongside improves performance, increasing ACC by 1.53% and AUC by 1.43%. This improvement highlights the benefits of incorporating contrastive learning between images and texts. The best performance is achieved by adding , reflecting the importance of fine-grained reconstruction in improving unimodal features extraction.
(2) Impact of model components. To assess the effectiveness of each proposed module, we conduct ablation experiments and report the results in Table 7(b). The configuration labeled “w/o KL, MEM” excludes both the knowledge (KL) and memory (MEM) injection modules, relying solely on visual and textual representations for predictions. The “w/o MEM” configuration removes only the memory module with PMM attention. The reported results confirm the critical role of both modules in AD diagnosis. Also, the use of our proposed KAG cross-attention further improves our performance.
(3) Impact of expert knowledge. The final ablation study examines the impact of expert knowledge, with the results presented in Table 8. This study evaluates the model’s performance with and without the integration of expert-provided knowledge. The findings demonstrate that incorporating expert knowledge significantly enhances diagnostic accuracy and AUC, underscoring its value in complementing the knowledge derived from LLM-generated general information.
Task | ACC(%) | AUC(%) | SEN(%) | SPE(%) |
(a) Loss Terms | ||||
90.264.41 | 87.376.05 | 94.422.61 | 80.3311.49 | |
90.753.02 | 87.422.33 | 95.482.21 | 79.388.78 | |
91.793.77 | 88.804.91 | 95.834.11 | 81.769.71 | |
HoloDx (ours) | 92.303.24 | 89.385.32 | 96.602.11 | 82.1610.83 |
(b) Module Terms | ||||
w/o KL, MEM | 89.233.40 | 86.235.79 | 92.930.49 | 79.549.48 |
w/o MEM | 91.283.07 | 87.944.94 | 94.923.00 | 80.9710.80 |
w/o KAG | 91.792.99 | 88.542.82 | 97.832.87 | 77.249.44 |
HoloDx (ours) | 92.303.24 | 89.385.32 | 96.602.11 | 82.1610.83 |
Type | ACC(%) | AUC(%) | SEN(%) | SPE(%) |
LLM | 91.996.53 | 90.817.54 | 83.6114.04 | 98.013.99 |
Expert | 92.024.98 | 92.644.41 | 96.007.99 | 89.299.97 |
LLM+Expert | 94.674.99 | 94.805.17 | 93.817.62 | 95.785.18 |

### v2

@@ Abstract
Accurate diagnosis of Alzheimer’s disease (AD) requires effectively integrating multimodal data and clinical expertise. However, existing methods often struggle to fully utilize multimodal information and lack structured mechanisms to incorporate dynamic domain knowledge. To address these limitations, we propose HoloDx, a knowledge- and data-driven framework that enhances AD diagnosis by aligning domain knowledge with multimodal clinical data. HoloDx incorporates a knowledge injection module with a knowledge-aware gated cross-attention, allowing the model to dynamically integrate domain-specific insights from both large language models (LLMs) and clinical expertise. A memory injection module with prototypical memory attention further enables consistency preservation across decision trajectories. Through synergistic operation of these components, HoloDx achieves enhanced interpretability while maintaining precise knowledge-data alignment. Evaluations on five AD datasets demonstrate that HoloDx outperforms state-of-the-art methods, achieving superior diagnostic accuracy and strong generalization across diverse cohorts. The source code is released at https://github.com/Qybc/HoloDx.
@@ 1 Introduction
Alzheimer’s disease (AD) is a progressive neurodegenerative disorder that leads to memory loss, cognitive decline, and ultimately, loss of independence. Accurate diagnosis of AD is crucial for implementing appropriate treatment strategies, planning for future care, and providing support to patients and their families. Recent advancements [1, 2, 3, 4] in passive immunotherapies, such as Lecanemab and Donanemab, have shown promise in targeting amyloid-beta plaques, a hallmark of AD pathology. However, diagnosing AD, especially in its early stages, remains challenging due to subtle clinical symptoms and overlap with normal aging processes. Traditional diagnostic approaches, often relying on unimodal data such as neuroimaging or clinical assessments, struggle to capture these nuanced changes [5, 6]. To address these limitations, multimodal learning has gained attention for its ability to integrate diverse data sources, such as neuroimaging, clinical records, and genetic information, offering a more holistic perspective on AD’s early pathological changes [7, 8, 9, 10]. Despite the promise of data-driven approaches, significant challenges remain: 1) data bias, which focuses on prevalent patterns while often neglecting rare and atypical cases, and 2) a lack of interpretability, as many models function as opaque black boxes, making it difficult to trace causal relationships and ensure reliability. These limitations hinder the consistency, robustness, and trustworthiness of existing computer-aided diagnosis (CAD) systems. Addressing these shortcomings requires a synergistic paradigm, as shown in Fig. 1, which integrates the strengths of multimodal data-driven models with domain knowledge, enhancing interpretability, reliability, and generalizability in the context of AD and its early diagnosis.
@@ 2.1 Multimodal CAD
The integration of diverse modalities of medical data has become a prominent approach in advancing medical diagnosis [14, 15, 16, 17], enabling machine learning models to identify disease patterns more effectively than relying on single modalities. For example, Kim et al. [18] develop a heterogeneous graph learning method to fuse multimodal medical data efficiently. Similarly, Zhou et al. [8] introduce the IRENE model, which employs Transformer-based architectures to combine modality-specific low-level embeddings for diagnostic purposes. Multimodal integration has also succeeded in diagnosing conditions such as breast cancer, where combining imaging and genomic information has improved diagnostic outcomes [19]. Chen et al. [7] propose a transformer-based framework for aligning and fusing multimodal medical data to classify Alzheimer’s disease. These efforts highlight the growing potential of multimodal CAD systems to improve diagnostic precision and reliability. However, these approaches often rely solely on data-driven techniques, which limits their ability to handle complex medical scenarios where domain expertise is crucial. Without integrating domain knowledge, such models may fail to resolve conflicts between modalities or interpret rare and ambiguous cases effectively. This gap underscores the need for combining multimodal data integration with domain knowledge to enhance interpretability, consistency, and diagnostic reliability.
@@ 2.2 Knowledge-driven Multimodal CAD
Our approach bridges this gap by dynamically enriching the knowledge bank with general medical insights derived from large language models (LLMs) and specific knowledge contributed by domain experts, tailored to the factors associated with each subject. This ensures that multimodal data is augmented with both broad and context-specific representations, enabling our method to provide more precise and interpretable predictions, effectively addressing key limitations in existing knowledge-driven frameworks.
@@ 3.1.1 Factor Extraction
Expert: The increase of pTau217 in the serum of AD patients is usually early and significant, and its diagnostic power is generally superior to that of Abeta42/40 and pTau181. Therefore, pTau217 holds high clinical value in the early diagnosis of Alzheimer’s disease. | |
@@ 3.1.2 Factor Knowledge Generation
Here, refers to a specific factor , as listed in Table 1. Examples include the Mini-Mental State Examination (MMSE) scores collected from cognitive tests. The generated responses offer comprehensive descriptions of each factor and its relevance to diagnosing AD, as presented in the LLM knowledge column of Table 2.
@@ 3.2 Multimodal and Knowledge Feature Extraction
Unimodal Encoder. We employ unimodal encoders to process both visual inputs and textual inputs . For the visual modality, a 3D Vision Transformer [6] is used to encode the input images into vector representations, generating an image embedding that includes a [CLS] token to represent the global image feature. For the textual modality, we use a Longformer Transformer [25], which maps long text representations into feature embeddings. Similarly, a [CLS] token is appended to the beginning of the text input to summarize the whole sentence representation. Each unimodal encoder consists of transformer blocks comprising self-attention (SA) layers and feed-forward networks (FFNs).
@@ 3.3 Knowledge Injection Module
These multimodal embeddings initially undergo processing through self-attention (SA) mechanisms to systematically model interactions between modalities:
@@
Finally, the full layer transformation integrates these components through residual learning:
@@
The final patient representation aggregates modality-specific information through the concatenation of [CLS] tokens:
where , , and correspond to the [CLS] tokens of visual, textual, and knowledge embeddings, respectively.
@@ 3.5.1 Alignment
We employ the image-text contrastive (ITC) loss [29] to align image features and text features generated by the image and text encoders. The ITC loss maximizes similarity for positive image-text pairs while suppressing negative pairs, implemented through normalized cross-entropy over all pairwise similarities. To enhance consistency between input data and domain knowledge in disease diagnosis, we propose the Knowledge-Data Contrastive loss (KDC). This mechanism maintains two dynamically updated queues: a Data queue storing concatenated multimodal features , and a Knowledge queue containing domain embeddings generated by a dedicated knowledge encoder. The KDC loss follows the contrastive paradigm of , pulling matched knowledge-data pairs closer while distancing mismatched pairs. We implement momentum encoders updated via exponential moving average (EMA) following BLIP[30] and ALBEF[31]. Specifically, the parameters of momentum image/text encoders () are updated as , where is the momentum coefficient and denotes the parameters of the corresponding online encoders. All momentum encoders operate without gradient backpropagation. This EMA-based strategy ensures feature consistency within the dynamically updated data and knowledge queues by decoupling momentum encoder optimization from the online model training. To prevent abrupt shifts in feature distribution, the queues are exclusively maintained using outputs from the momentum encoder.
@@ 4.1 Datasets and Settings
Following established clinical guidelines [42, 43, 44], we evaluate our model on two classification tasks. The first task distinguishes cognitively normal (NC) individuals from those with cognitive impairment (CI), including both mild cognitive impairment (MCI) and Alzheimer’s disease (AD). The second one focuses on differentiating NC from MCI, which is a critical stage for the early identification of AD. We split the ADNI, AIBL, and RENJI datasets subject-wise into training, validation, and test sets with proportions of 70%, 10%, and 20%, respectively. All structural MRI scans underwent standardized preprocessing, including skull stripping [45] to remove non-brain tissues and intensity normalization to harmonize voxel value distributions across scanners. Also, we perform zero-shot tests on OASIS and MIRIAD datasets to assess inter-center generalization.
@@
Baselines. To evaluate the performance of our model, we compare it with three groups of baselines: (1) Image-only group. This group includes two recent methods commonly used in image classification: 3D ResNet [33], and 3D ViT [6]. (2) Text-only group. We select three prominent text-only baselines: BERT [34], Roberta [35], and Longformer [25]. (3) Multimodal group. This group includes four recent transformer-based models that fuse multimodal information for classification: GIT [36], IRENE [8], AD-Trans [9], and Alifuse [7].
@@
Implementation details For pre-processing, each image is resized and cropped to , with a spacing of [1.0,1.0,1.0] and an image patch size of . The maximum input text length of textual input is set to 512. Following common settings in [32], we set all hidden feature sizes to 768 and the number of heads in multi-head attention (MHA) to 12. The number of layers in the unimodal encoder is set to 12, while the number of layers in the decoder is set to 6. The number of layers in the knowledge and the memory modules is set to 6. The iteration is 100. The values for are empirically set to 1, 1, and 1, respectively. Our model is trained on four NVIDIA RTX 3090 GPUs with a batch size of , using the AdamW optimizer with a learning rate of 2e-5. Evaluation metrics include diagnosis accuracy (ACC), Area Under Curve (AUC), sensitivity (SEN), and specificity (SPE).
Task | ACC | AUC | SEN | SPE |
(a) Loss Terms | ||||
89.23 | 81.70 | 96.42 | 74.33 | |
89.75 | 84.42 | 95.48 | 79.38 | |
91.09 | 84.80 | 95.83 | 81.76 | |
91.28 | 85.74 | 98.54 | 72.92 | |
91.92 | 88.23 | 97.14 | 79.31 | |
(b) Module Terms | ||||
Backbone | 89.34 | 83.18 | 96.43 | 71.93 |
Backbone + KL | 91.28 | 86.94 | 94.92 | 80.97 |
Backbone + KL( w/o KAG) + MEM | 88.72 | 80.87 | 99.28 | 62.45 |
Backbone + KL + MEM (ours) | 91.92 | 88.23 | 97.14 | 79.31 |
(c) Feature Terms | ||||
Image | 71.58 | 54.20 | 95.71 | 11.67 |
Clinical | 88.21 | 82.10 | 96.43 | 67.76 |
Image + Knowledge | 75.76 | 60.64 | 97.12 | 24.14 |
Clinical + Knowledge | 89.54 | 84.77 | 96.43 | 73.11 |
Image + Clinical + Knowledge | 91.92 | 88.23 | 97.14 | 79.31 |
(d) Cognitive Tests | ||||
w/o | 79.80 | 79.43 | 80.28 | 78.57 |
w/ | 91.92 | 88.23 | 97.14 | 79.31 |
Image | Non-image |
Age is 70 years old. Gender is Male. Psychiatric history is non-existent. Neurologic history is non-existent. Cardiovascular history is non-existent. Hepatic history is non-existent. Musculoskeletal history is non-existent. Endocrine-Metabolic history is non-existent. Gastrointestinal history is non-existent. Renal-Genitourinary history is non-existent. Smoking history is existent. Malignancy history is non-existent. MMSE score is 27. Delayed Recall score is 5. Immediate Recall score is 6. Thyroid Stim Hormone is 1.9. Vitamin B12 is 528.594. MCH is 31.5. MCHC is 33.3. Platelets is 268.0. RBC is 4.13. Hemoglobin is 13.0. WBC is 5.7. Serum Glucose is 82.874. Cholesterol (High Performance) is 193.325. |
Age is 81 years old. Gender is Male. Psychiatric history is non-existent. Neurologic history is non-existent. Cardiovascular history is existent. Hepatic history is non-existent. Musculoskeletal history is existent. Endocrine-Metabolic history is non-existent. Gastrointestinal history is non-existent. Renal-Genitourinary history is non-existent. Smoking history is non-existent. Malignancy history is non-existent. MMSE score is 27. Delayed Recall score is 3. Immediate Recall score is 7. Thyroid Stim Hormone is 5.19. Vitamin B12 is 433.718. MCH is 32.3. MCHC is 34.7. Platelets is 252.0. RBC is 3.52. Hemoglobin is 11.4. WBC is 4.3. Serum Glucose is 90.08. Cholesterol (High Performance) is 201.058. |
@@ 4.2.3 Ablation Studies
(1) Impact of loss terms. We evaluate the contribution of each loss, i.e., the classification loss , two contrastive losses and , and the combined restorative loss for images and texts. Incorporating all loss components is optimal for maximizing performance across all metrics, as reported in Table 7(a). Using only , the model performance experiences a reduction of 6.5% in AUC compared to the best-performing configuration. Adding and alongside improves performance, increasing AUC by 3.1%. This improvement highlights the benefits of incorporating contrastive learning between images and texts. The effectiveness of lies in its unique combination of knowledge anchoring, noise suppression, and convergence acceleration. By aligning multimodal features with domain-specific knowledge embeddings , establishes a stable semantic reference frame grounded in clinical ontologies and medical guidelines, preventing feature drift during modality fusion. Unlike traditional image-text contrastive objectives that only align paired modalities, this loss ensures all features remain consistent with evidence-based medical knowledge. Moreover, the knowledge queue serves as a semantic filter, suppressing non-discriminative variations commonly found in medical data and promoting attention to clinically relevant biomarkers. This effect is supported by our feature shapley analysis as shown in Fig. 3. Furthermore, accelerates early-stage training by providing strong semantic gradients; as shown in Fig. 3, models without this component require over more iterations to achieve comparable representation quality. These combined benefits make a catalytic component in our framework, significantly enhancing both robustness and learning efficiency. The highest diagnostic performance is achieved when the restoration loss is included in training. Removing this type of losses on images and texts lead to a notable performance decline (3.43% drop in AUC), confirming its essential role in preserving modality-specific features during cross-modal alignment. This highlights the importance of fine-grained reconstruction supervision in enhancing unimodal features quality and overall diagnostic effectiveness.
(2) Impact of model components. To evaluate the effectiveness of individual components, we perform ablation experiments by selectively activating the Knowledge Module (KL), Memory Module (MEM), and Knowledge-aware Gated Attention (KAG), as reported in Table 7(b). Notably, introducing the Knowledge Module alone improves the baseline AUC by 3.76%, and its combination with the Memory Module yields an additional 1.29% gain. The full model achieves a total performance improvement of 5.05% in AUC. Additionally, we replace the Knowledge-aware Gated Attention with standard additive attention. This substitution results in a significant performance drop, specifically 2.56% in accuracy and 6.07% in AUC, highlighting the critical role of gated attention in facilitating adaptive and effective feature fusion.
(3) Impact of feature terms. We conduct an ablation study to evaluate the contributions of individual and partial modality combinations. As shown in Table 7(c), five configurations were compared: (1) image features only, (2) clinical data only (including demographics, cognitive assessments, biospecimens, and genetic information), (3) image + knowledge (general knowledge and expert rules), (4) clinical + knowledge, and (5) the full model incorporating all three feature terms.
The results reveal two key observations. First, clinical data provides broader contextual information than imaging alone. While the image-only model achieved an AUC of 54.20% by capturing visual patterns, the clinical-only model reached 82.10% AUC by leveraging trends from cognitive tests and historical medical data. The substantial 27.90% performance gap indicates that clinical features encode systemic disease characteristics often missed by imaging. Second, the integration of clinical data with additional modalities yields the greatest performance gains. The “clinical + knowledge” combination achieved 84.77% AUC, while the full model integrating image, clinical, and knowledge modalities achieved the highest overall performance of 88.23% AUC, demonstrating the complementary value of multi-modal information fusion.
(4) Impact of cognitive scores and multimodal integration. To assess the potential dominance of cognitive scores such as MOCA and MMSE in the diagnostic process, we conduct an ablation experiment by excluding these features from the input. As shown in Table 7(e), removing MOCA and MMSE leads to a substantial performance drop: accuracy decreased by 12.12% and AUC dropped by 8.8% on the AIBL dataset. This result underscores the strong correlation between cognitive scores and AD diagnosis, aligning with their clinical use as standard screening tools.
To further ensure robust decision-making, our model jointly leverages imaging data and interpretable mechanisms, preventing over-reliance on a single modality. Table 8 presents representative cases in which patients with identical MMSE scores (i.e., MMSE = 27) receive different diagnostic outcomes based on structural MRI findings. For instance, patients showing pronounced hippocampal atrophy or ventricular enlargement are correctly identified as cognitively impaired (CI), while others with similar cognitive profiles but normal imaging features are classified as cognitively normal (CN). These results highlight the complementary nature of multimodal input, i.e., clinical data offer sensitive screening, while imaging data provide structural confirmation. As shown in Table 7(c), our multimodal design leads to a 6.13% and 34.04% AUC gain over clinical-only and image-only models, respectively. Such synergy reinforces the necessity of integrating both modalities for comprehensive AD assessment.
Type | AIBL | RENJI | ||||||
ACC | AUC | SEN | SPE | ACC | AUC | SEN | SPE | |
LLM | 90.25 | 84.67 | 97.73 | 71.59 | 91.20 | 90.67 | 83.99 | 98.14 |
Expert | 91.23 | 87.03 | 99.28 | 74.78 | 92.14 | 91.42 | 91.56 | 85.67 |
LLM+Expert | 91.92 | 88.23 | 97.14 | 79.31 | 94.28 | 95.60 | 93.92 | 92.69 |
(5) Impact of expert knowledge. We further investigate the contribution of expert knowledge in guiding the diagnostic model, with results reported in Table 9. This study compares model performance with and without the integration of expert-curated clinical knowledge. On AIBL, compared to LLM-based knowledge, using expert knowledge improves diagnostic AUC by 2.36%; then, combining it with LLM and expert knowledge further improves the AUC by 1.20%, demonstrating its added value over general-purpose LLM outputs. Similar improvements have been observed on the RENJI dataset. The strength of expert guidance lies in its ability to highlight clinically significant biomarkers that generic LLMs often overlook, as shown in Table 2 and 3. Overall, expert knowledge offers clinical precision and contextual fidelity, particularly through fine-grained and task-specific criteria such as education-adjusted MMSE cutoffs (as seen in Table 2). In contrast, the LLM contributes broader contextualization and comprehensive recall, linking findings to a wider range of medical knowledge, including rare phenotypes and underrepresented cases. That is, the expert helps anchor the reasoning process and correct potential hallucinations, while the LLM helps fill in gaps and broaden perspective. Together, they form a more balanced and complete diagnostic reasoning system.
(6) Impact of loss weighting coefficients. To examine the influence of loss weighting coefficients on model performance, we conduct an ablation study on the AIBL test set by varying the coefficients and associated with the alignment loss () and the restoration loss (). Similar to previous work on multi-loss optimization [1], we design five representative configurations to explore the trade-off between cross-modal alignment and modality-specific discriminability. As reported in Table 10, the balanced configuration (, ) achieves the highest performance (91.92% ACC, 88.23% AUC), demonstrating the effectiveness of jointly optimizing both loss terms.
We observe that over-emphasizing alignment (Case B: ) leads to a modest sensitivity gain (+0.79%) but at the cost of a 3.89% drop in specificity, indicating excessive feature homogenization. Conversely, prioritizing restoration (Case C: ) slightly improves AUC (87.03%) but degrades sensitivity, suggesting incomplete modality alignment. Removing either loss (Cases D and E) results in pronounced performance degradation, with the alignment-free setting (Case E) performing worst (89.23% ACC). These findings highlight the interdependence between alignment and restoration objectives. A balanced weighting scheme mitigates the risks of over-smoothing or under-alignment, supporting more stable and generalizable multimodal learning.
Case | ACC | AUC | SEN | SPE | ||
A | 1.0 | 1.0 | 91.92 | 88.23 | 97.14 | 79.31 |
B | 1.0 | 0.1 | 91.28 | 86.67 | 97.93 | 75.42 |
C | 0.1 | 1.0 | 90.76 | 87.03 | 96.28 | 74.77 |
D | 1.0 | 0.0 | 91.09 | 84.80 | 95.83 | 81.76 |
E | 0.0 | 1.0 | 89.23 | 86.28 | 93.65 | 78.91 |
Note: Optimal performance achieved with balanced weights (Case A). Excessive alignment focus (Case B) improves sensitivity at cost of specificity. Representation-only loss (Case E) shows weakest cross-modal alignment.